the pdf of a continuous random variable x is given by f(x)="\int c/" square root x where 0 <x<4 find the constant c and find the cumulative distribution function of x then compute P(x>1)


1
Expert's answer
2021-11-01T04:55:59-0400

On of the properties of pdf is: "\int_{-\infty}^{+\infty} f(x)dx = 1", where f(x) is the pdf.

In the given case: "\int_{-\infty}^{+\infty} {\frac c {\sqrt x}}dx = 1\to\int_0^4 {\frac c {\sqrt x}}dx = 1\to 2c\sqrt{4}-2c\sqrt{0}=1\to"

"\to 4c=1\to c = 0.25"

The cdf is equal to 0 when x≤0, equal to 1 when x≥4, and equal to "\int_0^x {\frac 1 {4\sqrt t}}dt" when x is 0<x<4

"\int_0^x {\frac 1 {4\sqrt t}}dt =0.25\int_0^x {\frac 1 {\sqrt t}}dt=0.25(2\sqrt{x}-2\sqrt{0})={\frac {\sqrt{x}} 2}"

"P(x>1) = 1-P(x\u22641)=1-0.5=0.5"